OpenStudy (anonymous):

A collection of quarters and nickels is worth $14.80. There are 88 coins in all. how many of each coin are there? quarters?? nickels??

OpenStudy (ybarrap):

There are n nickels and q quarters. The sum of the number of coins is 88: n+q=88 We also know that each nickel is worth .05 of a dollar and a quarter is .25 of a dollar. So the total value of these n+q coins is n*.05+q*.25=14.80 There are two equations with two unknowns. Easy to solve: n=88-q Then substitute n into second equation and solve for n. Then take this n and find q: q=88-n That's it! Hope this helps!

OpenStudy (anonymous):

@ybarrap That doesn't make sense

OpenStudy (fibonaccichick666):

ok, let's try it this way, your goal is to come up with two equations you can solve

OpenStudy (fibonaccichick666):

both equations need to include the number of nickels, let's call it n, and the number of quarters, lets call it q.

OpenStudy (fibonaccichick666):

So now, you are told you have 88 coins, right? How do they get that number? What do they do?

OpenStudy (fibonaccichick666):

@Austin122131

OpenStudy (ybarrap):

How many coins do you have @Austin122131 ?